# Differential Privacy for Secure Machine Learning in Healthcare IoT-Cloud Systems

### v1

@@ 1 Introduction
The Internet of Things (IoT) has become an integral part of the modern application ecosystem, transforming not only organisational and industrial operations but also our daily lives. In the healthcare sector, digitalisation has revolutionised patient care through enhanced data collection, personalised medicine and preventative treatment approaches. Wearable devices such as Fitbit trackers, smart glasses and insulin pumps now facilitate continuous monitoring of vital parameters including heart rate, blood sugar levels and sleep patterns. These Healthcare IoT devices require configuration with the user’s personal information such as age, gender and location. Healthcare Analytics is a branch of data analytics that harnesses this patient data to generate descriptive, diagnostic, predictive, and prescriptive insights, fundamentally supporting evidence-based medical decision-making. Moreover, through the application of AI/ML techniques, governmental health agencies and research institutions are developing socially beneficial applications such as epidemiological forecasting, disease progression modeling, healthcare resource allocation, and actuarial risk assessment.
@@
Medical data from IoT devices, imaging technologies (MRI, CT scans), test results, Electronic Health Records (EHR), medication monitoring, and environmental sensors are stored in Cloud systems, creating comprehensive repositories for medical history, diagnosis, research and analytics. This cloud infrastructure provides secure access to various stakeholders such as physicians, clinicians, paramedics, pharmaceutical companies, insurance firms, researchers, and analysts, enabling them to examine and analyse critical healthcare information [nikki]. However, this centralised data storage presents significant security vulnerabilities, with adversaries potentially exploiting patient information for blackmail, ransom demands, manipulation, generating artificial market demand, and targeted marketing campaigns [mn2021ccem]. The protection of patients’ data thus becomes critically important, as misuse can lead to numerous harms including psychological distress, compromised treatment through data corruption, and various forms of extortion.
@@
Healthcare expenditure in the USA reached nearly 20% of GDP in 2021, yet system security remains problematic, with 692 large healthcare data breaches reported between July 2021 and June 2022. To address these vulnerabilities, healthcare organisations must comply with regulatory frameworks such as the Health Insurance Portability and Accountability Act (HIPAA), established by the U.S. Congress in 1996 [kebi22bc]. HIPAA encompasses three primary components: the Security Rule, the Privacy Rule, and the Breach Notification Rule. These regulations mandate essential protective measures for Electronic Health Record systems, including access control, data encryption, and comprehensive audit trails [ocr23]. The HIPAA Privacy Rule, specifically, establishes national standards governing the security of health information during transfer, reception, processing and sharing [hipaa23]. Despite these regulatory safeguards, the persistence of data breaches underscores the need for more robust privacy-preserving techniques. Differential Privacy (DP) has emerged as a particularly effective approach to securing personally identifiable information, offering advantages over conventional methods through its mathematical framework that anonymises data by introducing calibrated noise to datasets [cdwork].
@@
Medical records contain various personal details of patients and are stored in cloud systems for use by multiple stakeholders. Suppressing explicit parameters such as name and address does not guarantee privacy of the patient because combinations of other fields such as postcode and date of birth can combine uniquely to re-identify the person.
A significant case illustrating privacy vulnerabilities in healthcare data occurred when Sweeney [sweeney] demonstrated how medical records could be re-identified despite anonymisation. In one instance, she successfully re-identified the medical records of the Governor of Massachusetts by executing merely three queries on an anonymised hospital dataset released by the Massachusetts Group Insurance Commission (GIC). The re-identification was facilitated by cross-referencing public information—news reports of the governor’s collapse on 18 May 1996 and subsequent hospitalisation—with voter registration records containing demographic identifiers (names, addresses, postcodes, birth dates, and gender). This case exemplifies how quasi-identifiers can be matched across datasets to compromise patient privacy. In contemporary healthcare systems, Machine Learning (ML) techniques are routinely employed to analyse medical data and address research questions [qayyum2020secure], yet these analytical capabilities simultaneously present heightened privacy risks if not properly safeguarded. By scripting fine-tuned sequences of ML queries, adversaries can steal personal information of patients from healthcare databases. DP can help to safeguard against such privacy issues.
The Differential Privacy guarantee states that in a database, adding or dropping a row does not affect the output of the analysis, and an adversary cannot tell if a particular person is present in the database or not. Consider, for instance, a health department investigating the causal relationship between smoking and cancer incidence. Potential participants may exhibit reluctance to engage in such research due to concerns regarding social stigma, legal implications, and possible insurance premium discrimination if their identities become traceable within the dataset. In this scenario, robust data privacy protections are essential to facilitate accurate analysis of the smoking-cancer relationship while ensuring participant anonymity. DP offers a methodological framework that effectively balances individual privacy preservation with data utility for research purposes. Although alternative approaches such as homomorphic encryption [wood2020homomorphic] and multi-party secure computing [knott2022cryptensecuremultipartycomputation] enable operations on protected data while maintaining confidentiality, these methods frequently present substantial computational complexity challenges [mn2023IJACSAuk]. The principal advantage of DP lies in its capacity to maintain data utility for statistical analyses while specifically obfuscating individual-level identifying information.
@@
The primary focus of this research is to enhance response time and security in healthcare data systems, with particular emphasis on protecting patient privacy. We employ DP techniques to safeguard patient data while maintaining its utility for analysis. While prior works have explored DP-ML integration [bin22a, kek20i] and blockchain-based healthcare systems [hong21pp, kebi22bc, zhan22it, moh23bs] independently, our work advances the state-of-the-art through three key innovations that distinguish it from existing approaches.
@@
Third, our integrated architecture uniquely combines three orthogonal security mechanisms, that is, DP for input privacy, blockchain for data integrity and provenance, and hierarchical access control across computing layers, into a unified framework optimized for healthcare emergency response scenarios. While prior work has explored pairwise combinations of these security mechanisms, our approach formally analyses the security guarantees of the complete integrated system under a comprehensive threat model encompassing inference attacks, data tampering, and consensus attacks. This holistic security analysis demonstrates that the three mechanisms provide complementary protections without introducing conflicting requirements or degrading system performance.
@@ 2.1 Machine Learning
ML is a branch of artificial intelligence that focuses on creating systems that learn patterns from data and improve automatically with experience. In healthcare, ML supports diverse applications including disease identification, predictive diagnostics, image analysis, epidemic control, treatment optimisation, genomic analysis, and surgical automation[qayyum2020secure]. These techniques enable researchers to analyse population-level data such as demographic patterns, symptom clusters, and disease correlations, classifying data based on features to derive insights for diagnosis, prediction, and analysis of diseases.
ML approaches can be categorised into two main types: supervised and unsupervised learning. Supervised learning utilises labelled data, where input features are paired with corresponding outputs. The algorithm learns to predict outputs for new inputs by recognising patterns from training data. This approach is commonly employed for classification, regression, and object detection tasks. Key supervised algorithms include Logistic Regression [logisticregression], Random Forests [randomforests], and Naive Bayes [naivebayes]. In regression problems, the output variable represents a continuous value (e.g., ’pounds’ or ’kilogrammes’), while classification problems involve categorical outputs (e.g., ’disease’ or ’no disease’). Unsupervised learning, conversely, operates on unlabelled data without predefined outputs. These algorithms identify inherent patterns and relationships within datasets, making them suitable for clustering, dimensionality reduction, and anomaly detection. K-Means [kmeans] represents a prominent unsupervised learning technique used for grouping similar data points into clusters based on their characteristics.
@@
(iii) Random Forest represents an ensemble learning technique that constructs multiple decision trees during training and aggregates their predictions, typically through majority voting. This approach mitigates overfitting and enhances generalisation, rendering it particularly suitable for both binary and multi-class classification challenges in medical contexts. Random Forest algorithms have demonstrated considerable utility in analysing complex medical datasets, including radiological images and electrocardiographic data, predicting disease progression, and identifying optimal therapeutic compounds based on molecular characteristics.
(iv) Naive Bayes classifiers apply Bayes’ theorem with the naive assumption of conditional independence among features given the class label. This probabilistic approach is computationally efficient and performs remarkably well in various text classification tasks despite its simplifying assumptions. In medical contexts, Naive Bayes algorithms effectively process symptom-based diagnostic classification and medical document categorisation. They have shown particular efficacy in diabetes detection when applied to clinical datasets, offering a balance between computational simplicity and predictive performance.
@@ 2.2 Differential Privacy
DP is a data sequestration technique that aims to safeguard individual privacy while enabling analysis of large datasets. This mathematical approach introduces controlled noise to data, making the outputs of queries differing by at most one record indistinguishable, thereby ensuring individual privacy [soria2017]. The DP guarantee for two datasets is given by:
@@ 2.2.1 Types of DP
DP can be classified along two primary dimensions. First, based on implementation architecture, into Centralised DP and Local DP. Second, based on methodology, into Pure DP, Query-Level DP, and Moment DP [deeplearningDP].
@@ 2.2.3 Underlying Concepts of DP
  • Probability to Fail/Probability of Error ($\delta$): The delta parameter accounts for the probability of privacy protection failure, specifically, the likelihood that a query might reveal an individual’s presence in the dataset. Such failures may occur $\delta\times n$ times, where $n$ represents the number of records [mar16y].
(ii) Centralised Versus Local DP: As discussed in Section 2.2.1, two principal implementation strategies exist for DP. Centralized DP employs a trusted curator who applies precisely calibrated noise to query results. This approach typically utilises Laplace or Gaussian noise mechanisms and is commonly referred to as the trusted curator model. Conversely, local DP operates without requiring a trusted intermediary, hence its designation as the untrusted curator model. In local DP implementations, data undergoes randomisation before curator access. A trusted entity may also employ local DP to simultaneously randomise all database records. Local DP algorithms frequently produce more heavily perturbed data, as noise is applied at the individual record level. This approach provides a particularly rigorous privacy standard with plausible deniability guarantees, establishing local DP as a leading methodology for privacy-preserving data collection and distribution, but at the cost of model utility.
(iii) Correlated Sensitivity: Whilst Global Sensitivity effectively measures the maximum number of correlated records, it fails to account for the degree of data correlation. The concept of Correlated Sensitivity addresses this limitation by quantifying the cascading impact on related records when a single record undergoes modification [toa19l]. This refinement proves particularly valuable when analysing datasets containing interdependent entries, a common characteristic of complex relational databases.
@@ 2.2.6 Applications of DP
(i) Healthcare: DP enables secure sharing of electronic health records and clinical trial data, allowing researchers to extract insights whilst maintaining patient confidentiality [sun2021pmrss, jiang2022attribute].
@@
(ii) Financial Services: Financial institutions employ DP for fraud detection, risk assessment, and customer behaviour modelling. This protects individuals’ financial information from exposure that could lead to identity theft or fraudulent activities [lin2020protecting].
@@
(iii) Social Media and Advertising: Platforms implement DP to analyse user behaviour for personalised experiences. Google has applied this approach in its advertising platforms to create user profiles without exposing individual behaviours, addressing growing privacy concerns in digital advertising [jiang2021applications, mobiledevmemo2021].
@@
(iv) Transportation: DP extracts insights from journey patterns and traffic flow data without revealing individual travel habits. This protection is essential as location data can reveal sensitive information about individuals’ routines and behaviours [cai2019differential].
@@
(v) Research and Academia: DP supports research by enabling examination of sensitive datasets while preserving participant anonymity. This balance is crucial for maintaining ethical standards in research whilst facilitating valuable data-driven discoveries [zhan2024preserving].
@@
(vi) Government and Policy-making: DP facilitates evidence-based policy development using large-scale datasets. This protection is necessary as government data often contains comprehensive citizen information that, if compromised, could lead to significant privacy violations [sar21c].
@@
(vii) Web Browsing Analysis: Google implemented DP techniques through their RAPPOR framework for Chrome browsing data to identify performance issues such as slow-loading web pages. This approach enables system improvements whilst protecting individual browsing patterns from surveillance or profiling [erlingsson2014rappor].
@@
(viii) Location-Based Services: Google Maps utilises DP to provide real-time traffic updates and location-based services from millions of users’ data. This implementation preserves user anonymity whilst delivering accurate collective insights about traffic conditions and business popularity [googledevelopers2019].
@@
(ix) Search Query Analysis: Google’s search engine employs DP to enhance recommendations by analysing queries and user interactions. This enables improved search results without compromising individual search histories, which may contain sensitive personal information [mobiledevmemo2021].
@@
(x) Public Health Monitoring: During the COVID-19 pandemic, Google released aggregated mobility reports using DP to protect individual identities. These reports assisted health authorities in understanding social distancing patterns whilst maintaining location privacy of individual users [aktay2020google].
2.2.7 Blockchain

### v2

@@ 1 Introduction
The Internet of Things (IoT) has become an integral part of the modern application ecosystem, transforming not only organizational and industrial operations but also our daily lives. In the healthcare sector, digitalization has revolutionised patient care through enhanced data collection, personalised medicine, and preventative treatment approaches. Wearable devices such as Fitbit trackers, smart glasses, and insulin pumps now facilitate continuous monitoring of vital parameters including heart rate, blood sugar levels, and sleep patterns. Users must configure these Healthcare IoT devices with personal information such as age, gender, and location. Healthcare Analytics leverages this patient data to generate descriptive, diagnostic, predictive, and prescriptive insights, thereby supporting evidence-based medical decision-making. Moreover, through AI/ML applications, governmental health agencies and research institutions are developing socially beneficial applications such as epidemiological forecasting, disease progression modeling, healthcare resource allocation, and actuarial risk assessment.
@@
Medical data from IoT devices, imaging technologies (MRI, CT scans), test results, Electronic Health Records (EHR), medication monitoring, and environmental sensors are stored in Cloud systems, creating comprehensive repositories for medical history, diagnosis, research, and analytics. This cloud infrastructure provides secure access to various stakeholders such as physicians, clinicians, paramedics, pharmaceutical companies, insurance firms, researchers, and analysts, enabling them to examine and analyse critical healthcare information [dang2019survey]. However, this centralised data storage presents significant security vulnerabilities. Adversaries may exploit patient information for blackmail, ransom demands, manipulation, creating artificial market demand, and targeted marketing campaigns [Mangala2021]. Therefore, protecting patient data is critically important, as misuse can lead to numerous harms including psychological distress, compromised treatment through data corruption, and various forms of extortion.
@@
Healthcare expenditure in the USA reached nearly 20% of GDP in 2021, yet system security remains problematic, with 692 large healthcare data breaches reported between July 2021 and June 2022. To address these vulnerabilities, healthcare organizations must comply with regulatory frameworks such as the Health Insurance Portability and Accountability Act (HIPAA), established by the U.S. Congress in 1996 [Azbeg2023]. HIPAA encompasses three primary components: the Security Rule, the Privacy Rule, and the Breach Notification Rule. These regulations mandate essential protective measures for Electronic Health Record systems, including access control, data encryption, and comprehensive audit trails [dohPrivacy]. The HIPAA Privacy Rule, in particular, establishes national standards governing the security of health information during transfer, reception, processing, and sharing [hipaa2023]. Despite these regulatory safeguards, persistent data breaches demonstrate the need for more robust privacy-preserving techniques. Differential Privacy (DP) has emerged as a particularly effective approach to securing personally identifiable information, offering advantages over conventional methods through its mathematical framework that anonymises data by introducing calibrated noise to datasets [Dwork2014].
@@
Medical records contain various personal details of patients and are stored in cloud systems for use by multiple stakeholders. Suppressing explicit parameters such as name and address does not guarantee patient privacy because combinations of other fields such as postcode and date of birth can uniquely re-identify individuals.
A significant case illustrating privacy vulnerabilities in healthcare data occurred when Sweeney [Sweeney] demonstrated how medical records could be re-identified despite anonymization. She successfully re-identified the medical records of the Governor of Massachusetts by executing merely three queries on an anonymized hospital dataset released by the Massachusetts Group Insurance Commission (GIC). Cross-referencing public information—news reports of the governor’s collapse on 18 May 1996 and subsequent hospitalization—with voter registration records containing demographic identifiers (names, addresses, postcodes, birth dates, and gender) facilitated the re-identification. This case exemplifies how quasi-identifiers can be matched across datasets to compromise patient privacy. In contemporary healthcare systems, Machine Learning (ML) techniques are routinely employed to analyse medical data and address research questions [Qayyum2021], yet these analytical capabilities simultaneously present heightened privacy risks if not properly safeguarded. By scripting fine-tuned sequences of ML queries, adversaries can steal personal information of patients from healthcare databases. DP can help safeguard against such privacy issues.
The Differential Privacy guarantee states that adding or removing a single row from a database does not significantly alter the output of the analysis, preventing an adversary from determining whether a particular person is present in the database or not. Consider, for instance, a health department investigating the causal relationship between smoking and cancer incidence. Potential participants may exhibit reluctance to engage in such research due to concerns regarding social stigma, legal implications, and possible insurance premium discrimination if their identities become traceable within the dataset. In this scenario, robust data privacy protections are essential to facilitate accurate analysis of the smoking-cancer relationship while ensuring participant anonymity. DP offers a methodological framework that effectively balances individual privacy preservation with data utility for research purposes. Although alternative approaches such as homomorphic encryption [Wood2021] and multi-party secure computing [Knott2021] enable operations on protected data while maintaining confidentiality, these methods frequently present substantial computational complexity challenges [MN2023]. The principal advantage of DP lies in its capacity to maintain data utility for statistical analyses while specifically obfuscating individual-level identifying information.
@@
This research enhances response time and security in healthcare data systems, with particular emphasis on protecting patient privacy. We employ DP techniques to safeguard patient data while maintaining its utility for analysis. While prior works have explored DP-ML integration [Jia2022, Gai2020], blockchain-based distributed ML [Rangwala2025blockchain], and blockchain-based healthcare systems [Wu2021, Azbeg2023, Zhang2022, Wazid2023] independently, our work advances the state-of-the-art through three key innovations that distinguish it from existing approaches.
@@
Third, our integrated architecture uniquely combines three orthogonal security mechanisms: DP for input privacy, blockchain for data integrity and provenance, and hierarchical access control across computing layers, into a unified framework optimized for healthcare emergency response scenarios. While prior work has explored pairwise combinations of these security mechanisms, our approach formally analyses the security guarantees of the complete integrated system under a comprehensive threat model encompassing inference attacks, data tampering, and consensus attacks. This holistic security analysis demonstrates that the three mechanisms provide complementary protections without introducing conflicting requirements or degrading system performance.
@@ 2.1 Machine Learning
ML is a branch of artificial intelligence that focuses on creating systems that learn patterns from data and improve automatically with experience. In healthcare, ML supports diverse applications including disease identification, predictive diagnostics, image analysis, epidemic control, treatment optimization, genomic analysis, and surgical automation[Qayyum2021]. These techniques enable researchers to analyse population-level data such as demographic patterns, symptom clusters, and disease correlations, classifying data based on features to derive insights for diagnosis, prediction, and analysis of diseases.
ML approaches can be categorised into two main types: supervised and unsupervised learning. Supervised learning utilises labelled data, where input features are paired with corresponding outputs. The algorithm learns to predict outputs for new inputs by recognising patterns from training data. This approach is commonly employed for classification, regression, and object detection tasks. Key supervised algorithms include Logistic Regression [Hosmer2013], Random Forests [Breiman2001], and Naive Bayes [Hand2001]. In regression problems, the output variable represents a continuous value (e.g., ’pounds’ or ’kilogrammes’), while classification problems involve categorical outputs (e.g., ’disease’ or ’no disease’). Unsupervised learning, conversely, operates on unlabelled data without predefined outputs. These algorithms identify inherent patterns and relationships within datasets, making them suitable for clustering, dimensionality reduction, and anomaly detection. K-Means [JB1967] represents a prominent unsupervised learning technique used for grouping similar data points into clusters based on their characteristics.
@@
(iii) Random Forest represents an ensemble learning technique that constructs multiple decision trees during training and aggregates their predictions, typically through majority voting. This approach mitigates overfitting and enhances generalization, rendering it particularly suitable for both binary and multi-class classification challenges in medical contexts. Random Forest algorithms have demonstrated considerable utility in analysing complex medical datasets, including radiological images and electrocardiographic data, predicting disease progression, and identifying optimal therapeutic compounds based on molecular characteristics.
(iv) Naive Bayes classifiers apply Bayes’ theorem with the naive assumption of conditional independence among features given the class label. This probabilistic approach is computationally efficient and performs remarkably well in various text classification tasks despite its simplifying assumptions. In medical contexts, Naive Bayes algorithms effectively process symptom-based diagnostic classification and medical document categorization. They have shown particular efficacy in diabetes detection when applied to clinical datasets, offering a balance between computational simplicity and predictive performance.
@@ 2.2 Differential Privacy
DP is a data sequestration technique that aims to safeguard individual privacy while enabling analysis of large datasets. This mathematical approach introduces controlled noise to data, making the outputs of queries differing by at most one record indistinguishable, thereby ensuring individual privacy [Soria-Comas2017]. The DP guarantee for two datasets is given by:
@@ 2.2.1 Types of DP
DP can be classified along two primary dimensions. First, based on implementation architecture, into Centralised DP and Local DP. Second, based on methodology, into Pure DP, Query-Level DP, and Moment DP [Abadi2016].
@@ 2.2.3 Underlying Concepts of DP
  • Probability to Fail/Probability of Error ($\delta$): The delta parameter accounts for the probability of privacy protection failure, specifically, the likelihood that a query might reveal an individual’s presence in the dataset. Such failures may occur $\delta\times n$ times, where $n$ represents the number of records [Abadi2016].
(ii) Centralised Versus Local DP: As discussed in Section 2.2.1, two principal implementation strategies exist for DP. Centralized DP employs a trusted curator who applies precisely calibrated noise to query results. This approach typically utilises Laplace or Gaussian noise mechanisms and is commonly referred to as the trusted curator model. Conversely, local DP operates without requiring a trusted intermediary, hence its designation as the untrusted curator model. In local DP implementations, data undergoes randomization before curator access. A trusted entity may also employ local DP to simultaneously randomise all database records. Local DP algorithms frequently produce more heavily perturbed data, as noise is applied at the individual record level. This approach provides a particularly rigorous privacy standard with plausible deniability guarantees, establishing local DP as a leading methodology for privacy-preserving data collection and distribution, but at the cost of model utility.
(iii) Correlated Sensitivity: Whilst Global Sensitivity effectively measures the maximum number of correlated records, it fails to account for the degree of data correlation. The concept of Correlated Sensitivity addresses this limitation by quantifying the cascading impact on related records when a single record undergoes modification [Zhang2020]. This refinement proves particularly valuable when analysing datasets containing interdependent entries, a common characteristic of complex relational databases.
@@ 2.2.6 Applications of DP
(i) Healthcare: DP enables secure sharing of electronic health records and clinical trial data, allowing researchers to extract insights whilst maintaining patient confidentiality [Sun2022, Jiang2022].
@@
(ii) Financial Services: Financial institutions employ DP for fraud detection, risk assessment, and customer behaviour modelling. This protects individuals’ financial information from exposure that could lead to identity theft or fraudulent activities [Lin2021].
@@
(iii) Social Media and Advertising: Platforms implement DP to analyse user behaviour for personalised experiences. Google has applied this approach in its advertising platforms to create user profiles without exposing individual behaviours, addressing growing privacy concerns in digital advertising [Jiang2023, Lindell2011].
@@
(iv) Transportation: DP extracts insights from journey patterns and traffic flow data without revealing individual travel habits. This protection is essential as location data can reveal sensitive information about individuals’ routines and behaviours [Cai2019].
@@
(v) Research and Academia: DP supports research by enabling examination of sensitive datasets while preserving participant anonymity. This balance is crucial for maintaining ethical standards in research whilst facilitating valuable data-driven discoveries [Zhan2024].
@@
(vi) Government and Policy-making: DP facilitates evidence-based policy development using large-scale datasets. This protection is necessary as government data often contains comprehensive citizen information that, if compromised, could lead to significant privacy violations [Saeidian2021].
@@
(vii) Web Browsing Analysis: Google implemented DP techniques through their RAPPOR framework for Chrome browsing data to identify performance issues such as slow-loading web pages. This approach enables system improvements whilst protecting individual browsing patterns from surveillance or profiling [Erlingsson2014].
@@
(viii) Location-Based Services: Google Maps utilises DP to provide real-time traffic updates and location-based services from millions of users’ data. This implementation preserves user anonymity whilst delivering accurate collective insights about traffic conditions and business popularity [google2019].
@@
(ix) Search Query Analysis: Google’s search engine employs DP to enhance recommendations by analysing queries and user interactions. This enables improved search results without compromising individual search histories, which may contain sensitive personal information [Lindell2011].
@@
(x) Public Health Monitoring: During the COVID-19 pandemic, Google released aggregated mobility reports using DP to protect individual identities. These reports assisted health authorities in understanding social distancing patterns whilst maintaining location privacy of individual users [Aktay2020].
2.3 Blockchain